OpenStudy (anonymous):

Pllllllease someone help :( FInd the general solution to dx/dt= 2x(1 − x/3) − xy dy/dt= −2y + 3xy

OpenStudy (anonymous):

if you can help me i'll love you forever .. i'm so stuck

OpenStudy (anonymous):

i think that all you do is find x=x(t) and y=y(t) for each but i dont know how!

OpenStudy (anonymous):

do you think you can help me :(

OpenStudy (anonymous):

It's simple without the last term xy, so I'm thinking!

OpenStudy (anonymous):

ok thank you .. i'm thinking that you solve and just keep one as a constant? i'm not sure.

OpenStudy (anonymous):

like the variable that your not solving for of course

OpenStudy (anonymous):

this assignemnt is due tm morning and if i dont do it i'm soo screwed but I'm seriously so confused!

OpenStudy (anonymous):

Is it familiar, as your lecture?

OpenStudy (anonymous):

not relaly to be honest

OpenStudy (anonymous):

Do you have any link for similar problem?

OpenStudy (anonymous):

i'm trying to figure it out using those .. did you get those links? it said i lost connection

OpenStudy (anonymous):

I did, my PC fluctuate all the time. I'm reading!

OpenStudy (anonymous):

ok, its like a combination of both it seems .. i'm not sure if i know what i'm doing..

OpenStudy (anonymous):

do you understand it at all?

OpenStudy (anonymous):

Nope! I'm getting help from other, see if it works (sighs)

OpenStudy (anonymous):

i just dont understand becuase its mixing the 2 ideas .. so i dont know if i'm supposed to try and combine it or, .. i suck

OpenStudy (anonymous):

2x(1 − x/3) − xy = 0 −2y + 3xy = 0 -> y ( -2 + 3x) = 0 => y = 0, x = 2/3

OpenStudy (anonymous):

Look like we solving system of equation by substitution:

OpenStudy (anonymous):

Replace y = 0 and x = 2/3 into (1)

OpenStudy (anonymous):

is that what were supposed to do? or are we only supposed to sub y=0 into the other one since in that equation were not solving for x at all?

OpenStudy (anonymous):

we have y = 0 and x = 2/3 to plug into: 2x(1 − x/3) − xy = 0 (1)

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

You figure it out?

OpenStudy (anonymous):

yes i did! thank you!

OpenStudy (anonymous):

y = 0 --> x = 0, x = 3

OpenStudy (anonymous):

2 intersection points ( 0, 0 ) and ( 3, 0)

OpenStudy (anonymous):

With x = 2/3 --> y = 14/9

OpenStudy (anonymous):

point ( 2/3, 14/9)

OpenStudy (anonymous):

So we have 3 solutions !

OpenStudy (anonymous):

Do you think we're fine here?

OpenStudy (anonymous):

how did you get (3,0) ? ok .. but since in each case your solving for one variable, you cant sub in both an x and a y into the first equation can you? like doesn't that not make sense?

OpenStudy (anonymous):

Of course :P One at the time x = 2/3 or y = 0

OpenStudy (anonymous):

So the whole concept is find intersection points

OpenStudy (anonymous):

but if its a dx/dt and then a dy/dt can we do that? ahhh ok ok!

OpenStudy (anonymous):

factor x, y out, then set = 0 --> x, y value from one equation.

OpenStudy (anonymous):

use it value to substitude into the other equation.

OpenStudy (anonymous):

ah ok ok!

OpenStudy (anonymous):

Happy :)

OpenStudy (anonymous):

Sarah ???

OpenStudy (anonymous):

ahah yes1 thank you!!!

OpenStudy (anonymous):

Finally ( relieve sighs)

OpenStudy (anonymous):

ahah sorry i was such a pain in the retrice

OpenStudy (anonymous):

You need to organize your notes

OpenStudy (anonymous):

Good thing is you have it to show me, so I figure out the whole idea is find intersection points

OpenStudy (anonymous):

Otherwise I don't even know what you want me to do !

OpenStudy (anonymous):

ahah i know i know .. i'm slacking a bit i'll admit, but totally appreciate you helping me .. want to help me with one more thing :) ? LAST thing i promise!

OpenStudy (anonymous):

I'm really exhausted today!

OpenStudy (anonymous):

Just post up to see if others better!

OpenStudy (anonymous):

PLEASE :) its just a little one From the differential equations determine (for x > 0, y > 0) where x and y are decreasing and where they are increasing.

OpenStudy (anonymous):

or if you even tell me how ..

OpenStudy (anonymous):

Then I need some sample as you did show me!

OpenStudy (anonymous):

its just for the DE above, the first ones we were working with!

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

dx/dt= 2x(1 − x/3) − xy

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Now what to do?

OpenStudy (anonymous):

i'm not sure how to find where its increasing and decreasing

OpenStudy (anonymous):

Let me search internet, look like need second direvative

OpenStudy (anonymous):

hhah awe, no dont worry about it! i'lll just leave it blank! go to sleep!

OpenStudy (anonymous):

It says, all we'll do is plug it the other value around (0,0) and (3,0) to get the signs

OpenStudy (anonymous):

suppose (1,0) or ( 0,1)

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

wish i understood this stuff, ok i'm just going to make something up.

OpenStudy (anonymous):

It's not difficult, just plug value around max,min.

OpenStudy (anonymous):

so its increasing if dy/dt<0 ? and decreasing if dx/dt<0?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

really! wow, i just guessed.

OpenStudy (anonymous):

You got it, At (1,0) , equation = 4/3 >0 => decreasing!

OpenStudy (anonymous):

awesome! thanks so much for all your help! i really appreciate it! would have died without ya! :)

OpenStudy (anonymous):

I did hear it from a guy today, I'm proud that I save couple of lives daily :P

OpenStudy (anonymous):

What time from your area now?

OpenStudy (anonymous):

its 12:33 am

OpenStudy (anonymous):

Oh same here!

OpenStudy (anonymous):

ahah then get some rest :) night!!

OpenStudy (anonymous):

What grade are you?